OpenStudy (anonymous):

solve: 7 / x + 2 – 4 / 2x -3

OpenStudy (anonymous):

does "solve" mean "subtract"?

OpenStudy (anonymous):

add or subtract and simplify the answer

OpenStudy (anonymous):

so you have \[\frac{7}{x+2}-\frac{4}{2x-3}\] \[=\frac{7(2x-3)-4(x+2)}{(x+2)(2x-3)}\] then some algebra in the numerator to remove parentheses and combine like terms

OpenStudy (anonymous):

o.k. is that the simplified answer?

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

= 14x - 21 - 4x - 8 -------------- (x + 2)(2x - 3} = 10x - 29 -------------- (x + 2)(2x - 3}